# Adaptive Task Allocation in Multi-Human Multi-Robot Teams under Team Heterogeneity and Dynamic Information Uncertainty

### v1

@@ I Introduction
Compounding this challenge is the highly dynamic nature of MH-MR team operations [6]. During task execution, human states, such as fatigue and engagement, can fluctuate over time, and robot operational conditions are also subject to change [7]. Concurrently, task specifications initially estimated may evolve as the operation progresses, introducing additional dynamics [8]. More crucially, in real-world scenarios, information about these dynamic states is often uncertain, influenced by inaccuracies in perception and communication latency [9]. For instance, human fatigue may not be measured with perfect accuracy, whether assessed subjectively [10] or objectively [11], and updates regarding robot status and task attributes can be delayed [12]. This necessitates adaptive task allocation methods that can respond to dynamic team states and changing conditions of tasks under uncertain information.
@@
  • We introduce ATA-HRL to hierarchically account for both inherent heterogeneity and dynamic operational states in MH-MR teams, optimized through a HRI structure.
@@ III-A Problem Formulation
    • $A_{\text{CTR}}:=\{A_{\text{C}}\times A_{\text{TR}}\}$ denotes a series of joint CTR actions at every step during operations, comprising two sub-components. $A_{\text{C}}=\sum_{t=1}^{T}a_{\text{C}}^{t}$ refers to the series of decisions on whether reallocation is necessary, determined by the reallocation condition policy $\pi_{\text{C}}:\mathcal{I}\times\mathcal{S}\times a^{t_{\text{Prev}}}_{\text{TR}}\mapsto a_{\text{C}}$ And $A_{\text{TR}}=\sum_{t=1}^{T}a_{\text{TR}}^{t}$ refers to the series of task reallocation decisions at each time step, determined by the task reallocation policy $\pi_{\text{TR}}:\mathcal{I}\times\mathcal{S}\times a^{t_{\text{Prev}}}_{\text{TR}}\mapsto a_{\text{TR}}$, which decides how tasks should be reassigned if reallocation is triggered. Here, $a^{t_{\text{Prev}}}_{\text{TR}}$ is the previous task reallocation decision, which is equivalent to $a_{\text{ITA}}^{t_{0}}$ at $t_{1}$.

### v2

@@ I Introduction
Compounding this challenge is the highly dynamic nature of MH-MR team operations [6, 7]. During task execution, human states, such as fatigue and engagement, can fluctuate over time, and robot operational conditions are also subject to change [8]. Concurrently, task specifications initially estimated may evolve as the operation progresses, introducing additional dynamics [9]. More crucially, in real-world scenarios, information about these dynamic states is often uncertain, influenced by inaccuracies in perception and communication latency [10]. For instance, human fatigue may not be measured with perfect accuracy, whether assessed subjectively [11] or objectively [12], and updates regarding robot status and task attributes can be delayed [13]. This necessitates adaptive task allocation methods that can respond to dynamic team states and changing conditions of tasks under uncertain information.
@@
  • We introduce ATA-HRL to hierarchically account for both inherent heterogeneity and dynamic operational states in MH-MR teams, optimized through a HRL structure.
@@ III-A Problem Formulation
    • $A_{\text{CTR}}:=\{A_{\text{C}}\times A_{\text{TR}}\}$ denotes a series of joint CTR actions at every step during operations, comprising two sub-components. $A_{\text{C}}=\sum_{t=1}^{T}a_{\text{C}}^{t}$ refers to the series of decisions on whether reallocation is necessary, determined by the reallocation condition policy $\pi_{\text{C}}:\mathcal{I}\times\mathcal{S}\times a^{t_{\text{Prev}}}_{\text{TR}}\mapsto a_{\text{C}}$. And $A_{\text{TR}}=\sum_{t=1}^{T}a_{\text{TR}}^{t}$ refers to the series of task reallocation decisions at each time step, determined by the task reallocation policy $\pi_{\text{TR}}:\mathcal{I}\times\mathcal{S}\times a^{t_{\text{Prev}}}_{\text{TR}}\mapsto a_{\text{TR}}$, which decides how tasks should be reassigned if reallocation is triggered. Here, $a^{t_{\text{Prev}}}_{\text{TR}}$ is the previous task reallocation decision, which is equivalent to $a_{\text{ITA}}^{t_{0}}$ at $t_{1}$.